OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (-6, 1) m = 6

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ -6}}\quad ,&{\color{blue}{ 1}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= 6 \\ \quad \\ y-y_1={\color{green} m}(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)

OpenStudy (anonymous):

my answer is y=6x-37. Is my answer correct

OpenStudy (jdoe0001):

\(\bf y-(1)={\color{green} 6}(x-(-6))\implies y-1=6(x+6)\implies y=6x+37\)

OpenStudy (anonymous):

Thanks. Can you help me with one more?

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

(10, 11) m = -9

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 10}}\quad ,&{\color{blue}{ 11}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -9 \\ \quad \\ y-(11)={\color{green} {-9}}(x-(10))\implies y=-9(x-10)+11\)